OpenStudy (anonymous):

what is 2x^+2x-4=0 ??? using factoring and quadratic formula ???

OpenStudy (anonymous):

(2x+4)(x-1)

OpenStudy (mathteacher1729):

1) http://www.wolframalpha.com 2) type "2x^+2x-4=0" 3) click "show steps" You should really learn how to do this yourself though... patrickjmt has lots of great videos and http://www.mathtv.com has some AWESOME videos under algebra then factoring Hope this helps. :)

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

(2x+4)(x-1) = 0 either 2x+4 = 0 or x-1 = 0 therefore x = 1 or x = -4/2 = -2

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

Jen try it yourself!! You know the method, you can do it :)